# First Author Affiliation / Address line 1 Affiliation / Address line 2

### v1

@@ Abstract
This paper introduces the UCFE: User-Centric Financial Expertise benchmark, an innovative framework designed to evaluate the ability of large language models (LLMs) to handle complex real-world financial tasks. UCFE benchmark adopts a hybrid approach that combines human expert evaluations with dynamic, task-specific interactions to simulate the complexities of evolving financial scenarios. Firstly, we conducted a user study involving 804 participants, collecting their feedback on financial tasks. Secondly, based on this feedback, we created our dataset that encompasses a wide range of user intents and interactions. This dataset serves as the foundation for benchmarking 12 LLM services using the LLM-as-Judge methodology. Our results show a significant alignment between benchmark scores and human preferences, with a Pearson correlation coefficient of 0.78, confirming the effectiveness of the UCFE dataset and our evaluation approach. UCFE benchmark not only reveals the potential of LLMs in the financial sector but also provides a robust framework for assessing their performance and user satisfaction.The benchmark dataset and evaluation code are available  [^1] https://anonymous.4open.science/r/UCFE-Benchmark-A278.
@@
Anonymous ACL submission
@@ 1 Introduction
Recent advances in large language models (LLMs) have expanded their potential applications in finance wu2023bloomberggpt; huang2023finbert; kim2024financial. Finance professionals are increasingly using LLMs to solve specialized financial tasks li2023cfgpt; xie2024openfinllmsopenmultimodallarge; yang2023fingptopensourcefinanciallarge; zhang2023xuanyuan20largechinese. Financial tasks are inherently complex, involving specialized context, financial terminologies, legal intricacies, and dynamic markets that involve information with high noise-to-signal ratiopagano1993financial; mullainathan2017machine; li2018more, which adds significant challenges for LLMs to address. Accurate analysis of financial information is crucial, as even minor ignorance in a signal or market information can lead to substantial financial losses tversky1981framing; thaler2008mental; mohamed2024size.
To be effective, LLMs need to swiftly adapt to fiscal policy changes, market fluctuation, extreme events, and global factors, identifying key signals within real-time data to manage volatility and mitigate risks gueta2024can; yadav2024generative. Financial markets can react instantly to news, making it crucial for LLMs to process information in near real-time by rapidly consolidating unstructured, real-time data from multiple sources nguyen2023generative; tong2024ploutos. Despite LLMs’ improving accuracy on tasks like sentiment analysis, market prediction, and risk assessment wimmer2023leveraging; lopez2023can; rizinski2024comparative, these models still face significant limitations, such as their reliance on static datasets and challenges in handling real-time data, which hinder their real-world applicability in dynamic financial contexts. Moreover, the evolving nature of financial regulation adds another layer of complexity, requiring LLMs to continuously update their knowledge to remain compliant and useful yao2024survey; he2024s. These limitations highlight the need for a more dynamic evaluation framework that assesses LLMs’ performance under real-time, evolving financial conditions, ensuring they can handle not only static tasks but also the unpredictable nature of real-world financial environments.
@@ 2.1 Financial Benchmark
FLARE xie2023pixiu evaluates models on five financial tasks 222sentiment analysis, news headline classification, named entity recognition, question answering, and stock movement prediction. Existing benchmarks zhang2023finevalchinesefinancialdomain; li2023chatgpt; yuan2024r primarily use multiple-choice questions to assess models’ domain knowledge, with questions sourced from real-world financial documents and publicly available financial reports or websites, covering a wide range of topics such as finance, economy, accounting, and certification. MMMU and MMMU-PRO yue2024mmmu; yue2024mmmu2 extend beyond traditional financial NLP tasks by incorporating multimodal inputs to better evaluate models in more complex financial tasks. Although these benchmarks have advanced the evaluation of financial language models, they predominantly consider structured NLP tasks with deterministic answers and rely heavily on multiple-choice questions or tasks with specific answers, such as sentiment analysis and named entity recognition. This limits their ability to assess generative capabilities, which are essential for simulating real-world financial applications krause2023large; koa2024learning.
@@ 2.2 User-Centric Framework
The implementation of user-centric models involves integrating users into core business processes to harness their creative potential, which has been successfully demonstrated by companies like LEGO, IBM, and Coloplast hienerth2011exploring; kwon2021enterprise. EUCA framework, a practical prototyping toolkit designed to make AI systems explainable to non-technical end-users, provides twelve end-user-friendly explanatory forms that do not require technical knowledge to bridge the gap between technical creators and non-technical users jin2021euca. In the financial domain, user-centric explainability is also crucial in algorithmic decision-making systems like robo-advisors naveed2022explainable; roveda2023human; pisoni2023responsible. Research has highlighted the importance of providing transparent and comprehensible explanations to users, which indicated that user trust and confidence in financial applications are positively correlated with the transparency and comprehensibility of the explanations provided how2020artificial; deo2021user; xu2024ai.
@@ 3 Background
Recent advancements in LLMs have demonstrated significant potential in addressing complex financial tasks. Numerous organizations are now actively training their own LLMs, aiming to enhance their performance by incorporating extensive domain-specific knowledge. For instance, FinGPT yang2023fingptopensourcefinanciallarge, which applies supervised fine-tuning to the LLaMA model, has shown notable improvements in financial tasks. While models like FinLLaMA xie2024openfinllmsopenmultimodallarge, through continued pretraining, have further advanced LLM performance across various metrics. These developments highlight the growing demand and potential for LLMs in the financial domain, both in academia and industry.
From a technical standpoint, LLMs have undoubtedly reduced costs and improved efficiency by quickly processing vast amounts of financial text data with commendable performance. However, the challenges of developing real-world financial applications extend beyond technical issues. These challenges include business requirements, industry-specific barriers, data privacy concerns, accountability, and ethical considerations nie2024survey; yao2024survey, along with a gap in understanding between LLMs, functioning as AI assistants, and the specific needs of financial experts.
@@ 5 UCFE Benchmark
In this section, we provide an overview of the technical details and evaluation pipeline of the UCFE benchmark. As shown in Figure 4, the evaluation starts by selecting finance-specific tasks (introduced in Section 4.3), where the model acts as an AI assistant. GPT-4o [^4] https://openai.com/index/gpt-4o-system-card/ simulates user interactions, generating dialogue data based on realistic behavior. Using LLMs to simulate user roles is common in recent research inaba2024largelanguagemodelsused. To minimize model bias, we established evaluation criteria (detailed in Section 5.2). Model outputs are then compared in pairs, with Claude-3.5-Sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet as the evaluator, following the common practice LLm-as-judge framework for evaluation liu2023gevalnlgevaluationusing. Each model’s performance is measured using Elo scores, which offer a relative comparison of capabilities. Finally, results are compared against human expert preferences to ensure evaluation robustness.
@@ 5.1 Evaluation Method
We use the Elo rating system for model evaluation, which is well-suited for comparing multiple models. This system is widely applied in competitive environments, such as match result prediction in sports like association football hvattum2010using; chiang2024chatbot. Its key advantages are:
@@ 5.2 Experimental Settings
To mitigate positional bias in LLM evaluations li2023splitmergealigningposition, we shuffled the input order during dialogue comparisons. To further minimize evaluator bias, such as misinformation or cognitive bias talboy2023challengingappearancemachineintelligence, we designed the evaluation prompts based on two key criteria:
@@
InternLM2.5-7B-Chat cai2024internlm2 General
Llama-3.1-70B-Instruct llama3modelcard General
@@
Llama3-XuanYuan3-70B-Chat zhang2023xuanyuan20largechinese Financial
Palmyra-Fin-70B-32k Palmyra-Fin-70B-32k Financial
Qwen2.5-14B-Instruct qwen2.5 General
@@ 5.3 Overall Results
To ensure robustness, we used two additional evaluators, Gemini-1.5-pro [^6] https://deepmind.google/technologies/gemini/pro/ and Deepseek-chat [^7] https://www.deepseek.com/. As seen in Figure 5, the final results across target models show minimal variation, suggesting that the LLM-as-evaluator framework introduces slight bias. Additionally, to examine potential length bias wei2024systematic, we analyzed dialogue lengths and turn counts across models (Figure 6). The analysis shows no significant correlation, indicating no observable bias.
@@ 5.5 Case Study
Previous research has demonstrated the significant influence of scaling laws on model performance kaplan2020scalinglawsneurallanguage; ruan2024observationalscalinglawspredictability. As shown in Figure 8 our results also show a similar trend. LLMs with larger parameters generally outperform smaller ones within the same series, also LLMs from the same backbone model have better results after being trained on financial corpus. However, Llama3.1 appears to be an outlier in this pattern. As shown in Figure 9, we illustrate a result of the Summarization task that highlights this phenomenon. In many of the data points, we observe that the 70B model tends to generate longer, more verbose outputs compared to smaller models. This aligns with the conclusions of chiang2024overreasoningredundantcalculationlarge, where larger models are prone to over-reasoning, generating lengthy and unnecessary responses to questions. In our evaluation framework, where clear scoring criteria have been established, these redundant outputs significantly lower the model’s performance.

### v2

@@ Abstract
This paper introduces the UCFE: User-Centric Financial Expertise benchmark, an innovative framework designed to evaluate the ability of large language models (LLMs) to handle complex real-world financial tasks. UCFE benchmark adopts a hybrid approach that combines human expert evaluations with dynamic, task-specific interactions to simulate the complexities of evolving financial scenarios. Firstly, we conducted a user study involving 804 participants, collecting their feedback on financial tasks. Secondly, based on this feedback, we created our dataset that encompasses a wide range of user intents and interactions. This dataset serves as the foundation for benchmarking 12 LLM services using the LLM-as-Judge methodology. Our results show a significant alignment between benchmark scores and human preferences, with a Pearson correlation coefficient of 0.78, confirming the effectiveness of the UCFE dataset and our evaluation approach. UCFE benchmark not only reveals the potential of LLMs in the financial sector but also provides a robust framework for assessing their performance and user satisfaction.The benchmark dataset and evaluation code are available  [^1] https://github.com/TobyYang7/UCFE-Benchmark.
@@
Yuzhe Yang1,thanks: Equal contribution., Yifei Zhang2,11footnotemark: 1, Yan Hu1, Yilin Guo1, Ruoli Gan1, Yueru He3, Mingcong Lei1, Xiao Zhang3, Haining Wang2, Qianqian Xie3,22footnotemark: 2, Jimin Huang3, Honghai Yu2,, Benyou Wang1,thanks: Corresponding author. 1The Chinese University of Hong Kong, Shenzhen, 2Nanjing University, 3The Fin AI yuzheyang@link.cuhk.edu.cn, yf_zhang@smail.nju.edu.cn,
@@ 1 Introduction
Recent advances in large language models (LLMs) have expanded their potential applications in finance Wu et al. (2023); Huang et al. (2023); Kim et al. (2024). Finance professionals are increasingly using LLMs to solve specialized financial tasks Li et al. (2023a); Xie et al. (2024); Yang et al. (2023); Zhang et al. (2023b). Financial tasks are inherently complex, involving specialized context, financial terminologies, legal intricacies, and dynamic markets that involve information with high noise-to-signal ratioPagano (1993); Mullainathan and Spiess (2017); Li et al. (2018), which adds significant challenges for LLMs to address. Accurate analysis of financial information is crucial, as even minor ignorance in a signal or market information can lead to substantial financial losses Tversky and Kahneman (1981); Thaler (2008); Mohamed et al. (2024).
To be effective, LLMs need to swiftly adapt to fiscal policy changes, market fluctuation, extreme events, and global factors, identifying key signals within real-time data to manage volatility and mitigate risks Gueta et al. (2024); Yadav et al. (2024). Financial markets can react instantly to news, making it crucial for LLMs to process information in near real-time by rapidly consolidating unstructured, real-time data from multiple sources Nguyen and Tulabandhula (2023); Tong et al. (2024). Despite LLMs’ improving accuracy on tasks like sentiment analysis, market prediction, and risk assessment Wimmer and Rekabsaz (2023); Lopez-Lira and Tang (2023); Rizinski et al. (2024), these models still face significant limitations, such as their reliance on static datasets and challenges in handling real-time data, which hinder their real-world applicability in dynamic financial contexts. Moreover, the evolving nature of financial regulation adds another layer of complexity, requiring LLMs to continuously update their knowledge to remain compliant and useful Yao et al. (2024); He et al. (2024). These limitations highlight the need for a more dynamic evaluation framework that assesses LLMs’ performance under real-time, evolving financial conditions, ensuring they can handle not only static tasks but also the unpredictable nature of real-world financial environments.
@@ 2.1 Financial Benchmark
FLARE Xie et al. (2023) evaluates models on five financial tasks 222sentiment analysis, news headline classification, named entity recognition, question answering, and stock movement prediction. Existing benchmarks Zhang et al. (2023a); Li et al. (2023b); Yuan et al. (2024) primarily use multiple-choice questions to assess models’ domain knowledge, with questions sourced from real-world financial documents and publicly available financial reports or websites, covering a wide range of topics such as finance, economy, accounting, and certification. MMMU and MMMU-PRO Yue et al. (2024a, b) extend beyond traditional financial NLP tasks by incorporating multimodal inputs to better evaluate models in more complex financial tasks. Although these benchmarks have advanced the evaluation of financial language models, they predominantly consider structured NLP tasks with deterministic answers and rely heavily on multiple-choice questions or tasks with specific answers, such as sentiment analysis and named entity recognition. This limits their ability to assess generative capabilities, which are essential for simulating real-world financial applications Krause (2023); Koa et al. (2024).
@@ 2.2 User-Centric Framework
The implementation of user-centric models involves integrating users into core business processes to harness their creative potential, which has been successfully demonstrated by companies like LEGO, IBM, and Coloplast Hienerth et al. (2011); Kwon et al. (2021). EUCA framework, a practical prototyping toolkit designed to make AI systems explainable to non-technical end-users, provides twelve end-user-friendly explanatory forms that do not require technical knowledge to bridge the gap between technical creators and non-technical users Jin et al. (2021). In the financial domain, user-centric explainability is also crucial in algorithmic decision-making systems like robo-advisors Naveed et al. (2022); Roveda et al. (2023); Pisoni and Díaz-Rodríguez (2023). Research has highlighted the importance of providing transparent and comprehensible explanations to users, which indicated that user trust and confidence in financial applications are positively correlated with the transparency and comprehensibility of the explanations provided How et al. (2020); Deo and Sontakke (2021); Xu et al. (2024).
@@ 3 Background
Recent advancements in LLMs have demonstrated significant potential in addressing complex financial tasks. Numerous organizations are now actively training their own LLMs, aiming to enhance their performance by incorporating extensive domain-specific knowledge. For instance, FinGPT Yang et al. (2023), which applies supervised fine-tuning to the LLaMA model, has shown notable improvements in financial tasks. While models like FinLLaMA Xie et al. (2024), through continued pretraining, have further advanced LLM performance across various metrics. These developments highlight the growing demand and potential for LLMs in the financial domain, both in academia and industry.
From a technical standpoint, LLMs have undoubtedly reduced costs and improved efficiency by quickly processing vast amounts of financial text data with commendable performance. However, the challenges of developing real-world financial applications extend beyond technical issues. These challenges include business requirements, industry-specific barriers, data privacy concerns, accountability, and ethical considerations Nie et al. (2024); Yao et al. (2024), along with a gap in understanding between LLMs, functioning as AI assistants, and the specific needs of financial experts.
@@ 5 UCFE Benchmark
In this section, we provide an overview of the technical details and evaluation pipeline of the UCFE benchmark. As shown in Figure 4, the evaluation starts by selecting finance-specific tasks (introduced in Section 4.3), where the model acts as an AI assistant. GPT-4o [^4] https://openai.com/index/gpt-4o-system-card/ simulates user interactions, generating dialogue data based on realistic behavior. Using LLMs to simulate user roles is common in recent research Inaba et al. (2024). To minimize model bias, we established evaluation criteria (detailed in Section 5.2). Model outputs are then compared in pairs, with Claude-3.5-Sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet as the evaluator, following the common practice LLm-as-judge framework for evaluation Liu et al. (2023). Each model’s performance is measured using Elo scores, which offer a relative comparison of capabilities. Finally, results are compared against human expert preferences to ensure evaluation robustness.
@@ 5.1 Evaluation Method
We use the Elo rating system for model evaluation, which is well-suited for comparing multiple models. This system is widely applied in competitive environments, such as match result prediction in sports like association football Hvattum and Arntzen (2010); Chiang et al. (2024). Its key advantages are:
@@ 5.2 Experimental Settings
To mitigate positional bias in LLM evaluations Li et al. (2023c), we shuffled the input order during dialogue comparisons. To further minimize evaluator bias, such as misinformation or cognitive bias Talboy and Fuller (2023), we designed the evaluation prompts based on two key criteria:
@@
InternLM2.5-7B-Chat Cai et al. (2024) General
Llama-3.1-70B-Instruct AI@Meta (2024) General
@@
Llama3-XuanYuan3-70B-Chat Zhang et al. (2023b) Financial
Palmyra-Fin-70B-32k team (2024) Financial
Qwen2.5-14B-Instruct Team (2024) General
@@ 5.3 Overall Results
To ensure robustness, we used two additional evaluators, Gemini-1.5-pro [^6] https://deepmind.google/technologies/gemini/pro/ and Deepseek-chat [^7] https://www.deepseek.com/. As seen in Figure 5, the final results across target models show minimal variation, suggesting that the LLM-as-evaluator framework introduces slight bias. Additionally, to examine potential length bias Wei et al. (2024), we analyzed dialogue lengths and turn counts across models (Figure 6). The analysis shows no significant correlation, indicating no observable bias.
@@ 5.5 Case Study
Previous research has demonstrated the significant influence of scaling laws on model performance Kaplan et al. (2020); Ruan et al. (2024). As shown in Figure 8 our results also show a similar trend. LLMs with larger parameters generally outperform smaller ones within the same series, also LLMs from the same backbone model have better results after being trained on financial corpus. However, Llama3.1 appears to be an outlier in this pattern. As shown in Figure 9, we illustrate a result of the Summarization task that highlights this phenomenon. In many of the data points, we observe that the 70B model tends to generate longer, more verbose outputs compared to smaller models. This aligns with the conclusions of Chiang and yi Lee (2024), where larger models are prone to over-reasoning, generating lengthy and unnecessary responses to questions. In our evaluation framework, where clear scoring criteria have been established, these redundant outputs significantly lower the model’s performance.